# Decouple and Decompose: Scaling Resource Allocation with DeDe

### v2

@@ 7 Evaluation
    How do DeDe and state-of-the-art approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
@@
    What is the contribution of each design in DeDe? (§7.3)
We focus on four approaches throughout our evaluation:
  • Exact Sol.  Exact Sol solves the original resource allocation problem using commercial optimization solvers. We reuse the implementation from POP, which employs Gurobi [22], cvxpy [17, 4], and CPLEX [15] for traffic engineering, cluster engineering, and load balancing, respectively.
  • POP.  POP-$k$ [43] randomly splits the problem into $k$ smaller subproblems, applies commercial solvers in each, and coalesces the resulting $k$ sub-allocations into a global allocation. However, POP only simulates the parallel execution by solving subproblems sequentially, collecting the solving time of each, and calculating the parallel solving time mathematically.
  • DeDe.  DeDe is our truly parallel implementation, with details described in Section 6.
  • DeDe.  To facilitate a fair comparison with POP, we create a variant of DeDe, denoted as DeDe, following POP-style simulation where subproblems are solved sequentially and a parallel solving time is calculated mathematically.
We also consider domain-specific approaches apart from these four approaches. In traffic engineering, we evaluate a demand-pinning approach (“Pinning” for short) similar to a prior work [41], with the top 10% of demands allocated using optimization engines and the remaining assigned to the shortest paths. We also evaluate Teal [63], a learning-accelerated traffic engineering approach with massive parallelism achieved on the GPU. In cluster scheduling and load balancing, we evaluate the greedy heuristic algorithms, Gandiva [61] and E-Store [52], respectively.
We compare these approaches based on their allocation quality and computation times. The computation time is measured on 64 CPU cores (2$\times$ Intel Xeon Gold 6142); an additional GPU (Nvidia Titan RTX) is made available for Teal.
@@
In this section, we compare DeDe against state-of-the-art approaches on resource allocation problems from Section 5.
@@ 7.1.1 Traffic Engineering
Figure 3: Comparison for the problem of maximizing total flow in traffic engineering.
@@
We adopt the largest network in Teal’s dataset for traffic engineering problems. This 1,739-node network is adapted from an AS-level internet topology [12] for WAN purposes, with the traffic data collected from a global cloud provider and then mapped into the large network.
@@
Maximize total flow.  Figure 3 compares DeDe to state-of-the-art methods in maximizing total flow. DeDe achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, demonstrating a superior trade-off between satisfied demand and computation time compared to Exact Sol, POP, and Pinning. While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer at 1658 seconds. POP-16 and POP-64 improve parallelism by dividing the workload into more subproblems, reducing computation times to 456 seconds and 380 seconds, respectively. Note that the speedup of POP-64 is limited compared to POP-16 as the number of cores per subproblem drops from 4 to 1. This increased parallelism also comes at a cost, with satisfied demand decreasing to 87.4% for POP-16 and 81.6% for POP-64. Pinning focuses on top demands to reduce problem size but remains constrained by the sequential nature of its optimization solvers. As a result, it achieves a modest speedup, requiring 149 seconds to reach 87.3% satisfied demand. Teal, a learning-accelerated approach, leverages massive GPU parallelism to achieve 89% satisfied demand in just 1 second. However, it relies on a highly customized machine-learning framework, demanding significant human effort for design and training.
DeDe, representing the same algorithm under ideal conditions as POP, achieves a satisfied demand that is 8% higher than DeDe within 10 seconds in Figure 3 and consistently delivers superior allocation quality across the other figures. The reason is that DeDe executes each iteration more quickly, allowing it to complete more iterations within a given time, resulting in higher allocation quality. The difference in iteration time between DeDe and DeDe arises from three factors, as discussed in more detail in § 7.3: (1) DeDe measures the end-to-end time, including problem-solving, compilation, solution unpacking, and other associated tasks, while DeDe only takes the problem-solving time into account. (2) DeDe’s parallel implementation (§ 6) incurs overhead from cache contention, which slows down the problem-solving process. (3) In DeDe, each subproblem is statically pre-assigned to one of the processes, making it susceptible to stragglers. In contrast, DeDe assumes perfect dynamic scheduling.
@@
Figure 4: Comparison for the problem of minimizing max link utilization in traffic engineering.
@@
Min-max link utilization.  Figure 4 compares DeDe with other approaches in minimizing maximum link utilization. DeDe achieves a low maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds. Exact Sol reaches 1.63 with an average computation time of 35 seconds, and it is faster than itself in maximizing total flow because minimizing max link utilization is "easier" and requires fewer iterations. Pinning takes 42.5 seconds, making it slower than DeDe and slightly slower than Exact Sol, as it requires additional time to rebuild the problem when the top 10% of node pairs change. POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 with 9, 2, and 0.4 seconds on average, respectively. Although POP-16 is faster than DeDe, a fairer comparison between POP-16 and DeDe, both simulating parallelism, reveals that DeDe achieves a comparable computation time to POP-16. The domain-specific solution Teal takes only 0.3 seconds with massive parallelism from GPU with a maximum link utilization of 1.74.
7.1.2 Cluster Scheduling
We evaluate a large cluster scheduling problem of allocating jobs to 5000 GPU/CPUs of 196 different types. We collect real throughput from two computing benchmarks [25, 35]. The number of GPU/CPUs for each type and the required number in each job is generated based on the distribution observed in traces from real GPU-sharing clusters [58]. To reduce simulation time, we capture snapshots through the scheduling process rather than executing the entire simulation from start to finish.
Figure 5: Comparison for the min-max allocation problem in cluster scheduling.
Max-min allocation.  Figure 5 compares DeDe against other approaches with the objective of max-min allocation. DeDe achieves a near-optimal max-min allocation of 0.97 within 5 seconds and 0.98 within 10 seconds, while DeDe reaches 0.98 within 2 seconds. Due to their sequential nature, Exact Sol and Gandiva encounter a slow computation time of 86 and 22 seconds. The greedy heuristic Gandiva even experiences a 0.56 low max-min allocation. POP-4 deliverers a comparable time-accuracy tradeoff as DeDe but is 3.4$\times$ slower than DeDe, which is simulated in the same way as POP did for a fair comparison. A faster POP variant, POP-16, exhibits a lower max-min allocation of 0.94, as it becomes more difficult for jobs to assess the ideal GPU/CPU among the limited selection in smaller subproblems.
Figure 6: Comparison for the problem of maximizing proportional fairness in cluster scheduling.
@@ 7.1.2 Cluster Scheduling
Figure 8: Comparison for the problem of minimizing shard movements in load balancing.
Proportional fairness.  Figure 6 compares DeDe against other approaches in maximizing proportional fairness. As the optimization problem becomes more difficult with the convex logarithm function in the objective, both DeDe and DeDe achieve a greater speedup than POP by thoroughly breaking down the problem into smaller and easier pieces. DeDe and DeDe reach a proportional fairness score of over 1,000 in 300 seconds and 100 seconds, respectively. In contrast, POP-4 takes 1,825 seconds to reach the same score, while POP-16 and POP-64 remain below 900 despite their shorter computation times. Exact Sol, though accurate, is prohibitively slow, requiring 12,432 seconds.
7.1.3 Load Balancing
We evaluate a load balancing problem of assigning 2048 data shards to 256 servers. In each round, we compute a new allocation of shards to servers given the new query loads [43].
Minimize shard movements.  Figure 8 compares DeDe with state-of-the-art approaches in minimizing shard movements. DeDe achieves an average of 24.1 shard movements in 15 seconds, slightly better than POP-4’s 25.4 shard movements within 16.9 seconds. In faster scenarios, DeDe achieves 35.1 shard movements in 2 seconds, while POP-4 achieves slightly fewer movements of 32.2 in 1.7 seconds. In a more fair comparison with POP, DeDe consistently costs 15–18% fewer shard movements than POP given a similar time limit. Exact Sol achieves the minimum shard movements of 13.2 but is significantly slower, taking an average of 1,572 seconds due to the NP-hardness of solving the mixed-integer linear programming problem inherent to load balancing. On the other hand, the greedy heuristic represents the opposite end of the time-accuracy trade-off, completing in just 25 milliseconds but requiring 50 shard movements.
@@ 7.1.3 Load Balancing
(c) Comparison to more optimization methods.
@@ 7.2 Robustness
We evaluate the robustness of DeDe and other baselines against changes in problem granularity, temporal fluctuations, spatial redistribution, and link failures. In all these evaluations, DeDe is run for 30 seconds. To mitigate the impact of the changing resources and demands, we normalize the satisfied demand by the optimal satisfied demand from Exact Sol.
Changes of problem granularity.  The granular problem required by POP, where resources are mutually interchangeable, is impractical in reality. In traffic engineering, for example, demands are constrained to capacity resources on pre-configured paths. Similarly, in cluster scheduling, 33% of GPU tasks in production clusters are limited to specific GPU types [58]. Using traffic engineering as an example, we evaluate all approaches across varying levels of problem granularity, as depicted in Figure 7(a). To quantify resource interchangeability, we use the mean edge betweenness centrality [10, 11], which measures the average percentage of demands served by a given edge.
@@
Temporal fluctuation.  We introduce temporal fluctuations [63] to the traffic matrices for robustness assessment. For each demand, we calculate the variance $\sigma^{2}$ in its changes between consecutive time slots and create a new normal distribution $N(0,k\sigma^{2})$. We choose the values of $k$ to be 2, 5, 10, and 20. Next, we randomly draw a sample from the newly created normal distribution and add it to each demand in every time slot, simulating temporal fluctuations.
@@
Spatial redistribution.  To evaluate robustness against spatial redistribution, we reassign traffic demands among node pairs. Specifically, we adjust the top 10% of demands, which initially account for 88.4% of the total volume, so that they take up 80%, 60%, 40%, and 20% of the total volume instead. Figure 7(c) shows that DeDe consistently fulfills the highest demand across all spatial distributions. Pinning’s performance decreases by 8.3%, as its strategy of focusing only on top demands relies on the presence of a heavy-tailed demand distribution. Teal has a slight decrease up to 2.4% due to unseen demand distribution in training. This is a relatively smaller decrease than the temporal fluctuation, as Teal can still handle simple, evenly distributed traffic demands. POP has a slight increase of 1.4% as demands are more granular, where no demand requires a significant fraction of the total resources.
Link failures.  In practice, large-scale inter-datacenter link failures are exceedingly rare. Even in the event of fiber link failures, they typically do not result in capacity loss on an inter-datacenter link, unless caused by a fiber cut [51]. As a stress test, however, we assess extreme failure scenarios described in previous work [67] and deliberately introduce 50, 100, and 200 link failures. Figure 10 illustrates the offline satisfied demand following these link failures, and we observe a consistent trend across different link failure scenarios across all approaches. Specifically, all approaches drop 0.6–1.1% under 50 link failures, 1.5–2.1% under 100 link failures, and 3–4.2% under 200 link failures. The reason is that these link failures constitute only a minor portion of the total 8,558 links, allowing all approaches to exhibit robustness given adequate time for re-computations.
@@ 7.3 Micro-benchmarks
We conduct micro-benchmarks to evaluate the impact of DeDe’s design components, with all evaluations performed in maximizing total flow in traffic engineering.
@@
Speedup when varying CPU cores.  Figure 9(a) evaluates the speedup of DeDe, DeDe, and Exact Sol with 1, 4, 16, and 64 CPU cores. DeDe achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores. DeDe also achieves an almost linear speedup when the number of CPU cores grows from 1 to 16. However, this linearity diminishes with 64 cores, where DeDe only achieves 18.2$\times$ speedup. This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in Section 7.1. In contrast to the linear speedup of DeDe and DeDe, the speedup of Exact Sol is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores. This is because of commercial solvers’ sequential nature, e.g., the simplex method [45] in linear programming requires thousands to millions of small steps to reach the optimal solution.
Convergence rate.  Both Teal and DeDe employ ADMM for parallel optimization, but their approaches differ. Teal’s ADMM adapts its computations based on matrix multiplication to leverage GPU calculation capabilities. Figure 9(b) illustrates the convergence rates of Teal and DeDe under various initialization. DeDe takes 30 seconds to surpass 90% satisfied demand. However, when initialized with Teal’s coarse neural network solution, DeDe achieves this within 20 seconds. Conversely, with naive initialization, DeDe requires more than 60 seconds. Teal, benefiting from the massive parallelism of GPUs, reaches 90% satisfied demand within 10 seconds. However, its adapted ADMM limits its performance, capping its satisfied demand at 91% at 60 seconds compared to DeDe’s 92%.
Comparison to more optimization methods.  Figure 9(c) investigates more constrained optimization methods other than DeDe’s ADMM method in maximizing the total flow in traffic engineering. Section 7.1 has covered the computation time of the optimal Exact sol, the suboptimal POP-64, and the near-optimal DeDe. As DeDe replaces constraints with additional terms in the objective function, we compare similar methods, including the penalty method and the augmented Lagrangian method. The augmented Lagrangian method is the same as in DeDe’s ADMM but without decoupling and decomposition, and thus it is over 3$\times$ slower than DeDe in reaching a satisfied demand of over 90%. The penalty method, which gradually increases penalty parameters from 0 to infinity to enforce the converted constraints in the objective function, converges slowly and is over 30$\times$ slower than DeDe in reaching a satisfied demand of over 90%.

### v3

@@ 7 Evaluation
    How do DeDe and baseline approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
@@
    What is the individual contribution of each design component in DeDe? (§7.3)
Our evaluation compares four approaches:
  • Exact sol.  This baseline solves the original resource allocation problem using commercial solvers. We adopt the implementation from POP, which utilizes cvxpy [15, 3], Gurobi [22], and CPLEX [25] for cluster engineering, traffic engineering, and load balancing, respectively.
  • POP.  POP-$k$ [44] randomly splits the problem into $k$ smaller subproblems, applies commercial solvers to each, and coalesces the resulting $k$ sub-allocations into a global allocation. However, POP only simulates the parallel execution by solving subproblems sequentially and calculating the parallel solving time mathematically.
  • DeDe.  DeDe is our truly parallel implementation (§6). By default, the solution from the previous optimization interval is used to warm-start the subsequent interval.
  • DeDe.  For a fair comparison with POP, we create a variant of DeDe, denoted as DeDe, following POP’s simulation methodology. DeDe solves subproblems sequentially and estimates the parallel solving time mathematically.
We also consider domain-specific approaches apart from these four approaches. In cluster scheduling and load balancing, we evaluate the greedy heuristic algorithms, Gandiva [63] and E-Store [53], respectively. In traffic engineering, we evaluate a demand-pinning approach (“Pinning”) similar to a prior work [42], where the top 10% of demands are allocated using optimization engines and the rest are assigned to shortest paths. We also evaluate Teal [65], a learning-accelerated traffic engineering approach that achieves massive parallelism using a GPU.
All approaches are evaluated in terms of allocation quality and computation time. The computation time is measured using 64 CPU cores (2$\times$ Intel Xeon Gold 6142), with an additional GPU (Nvidia Titan RTX) available for Teal.
@@
In this section, we compare DeDe against state-of-the-art approaches on resource allocation problems from §5.
7.1.1 Cluster Scheduling
We stress-test DeDe’s scalability in a large, heterogeneous cluster scheduling environment, where jobs are allocated in a time-sliced manner across 456 different types of compute resources, e.g., GPU/CPU instances that vary in vendor, generation, memory capacity, and other specifications. The quantity of each resource type is randomly drawn from $\{8,16,24,\ldots,64\}$, resulting in a total of 16,520 instances. We utilize the simulator in Gavel [45] and model job arrivals as a Poisson process with an average inter-arrival of 100 seconds. We update job allocation decisions every 6 minutes, repeating this process for 200 scheduling rounds (20 hours). For each job, the number of requested resource instances per type is drawn from $\{1,2,4,8,16,32\}$, and following prior work [59], 33% of the jobs are restricted to specific resource types. Job throughput is derived from relevant benchmarks [26, 36, 19]. Please refer to Appendix A for more details on this setup.
Figure 4: Results of the cluster scheduling variant that maximizes the minimum job throughput (max-min allocation).
Max-min allocation.  Figure 4 compares DeDe against other approaches in terms of max-min allocation. DeDe quickly attains a near-optimal max-min allocation of 0.94 in just 3 seconds and 0.99 within 10 seconds, demonstrating a better balance between max-min allocation and computation time compared with Exact sol., Gandiva, and POP. In contrast, Exact sol. experiences a significantly slower computation time of 156 seconds due to its sequential nature. The greedy heuristic, Gandiva, performs poorly, achieving only a max-min allocation of 0.43 (in 1.4 seconds). POP-4 provides a comparable tradeoff between time and accuracy to DeDe, but is 1.6$\times$ slower. The faster variant, POP-16, splits the problem into 16 smaller subproblems and achieves a lower max-min allocation of 0.9 (in 3.1 seconds), since smaller subproblems limit the ability to select optimal GPU/CPU assignments.
DeDe, representing the same algorithm under idealized conditions similar to POP’s simulation, achieves a normalized max-min allocation of 0.99 and is 3.3$\times$ faster than DeDe in Figure 4. The reason is that DeDe executes each iteration more quickly and completes more iterations within a given time. The discrepancy in iteration time between DeDe and DeDe arises from three factors: (1) DeDe measures the end-to-end time, including problem solving, compilation, solution unpacking, and other associated tasks, while DeDe only accounts for the core solving time; (2) DeDe’s parallel implementation incurs overhead from cache contention, which slows computation; (3) In DeDe, each subproblem is statically pre-assigned to one of the processes, making it susceptible to straggler delays. In contrast, DeDe assumes perfect dynamic scheduling.
Figure 5: Results of the cluster scheduling variant that maximizes proportional fairness among job throughputs.
Proportional fairness.  Figure 5 compares DeDe against other approaches in maximizing proportional fairness. This problem is particularly challenging due to the logarithmic form of the objective function. Exact sol., which uses the SCS solver in cvxpy, fails to reach optimality even after 5 hours of computation. As a result, several methods report normalized fairness scores that exceed 1 when normalized relative to Exact sol. Both DeDe and DeDe achieve greater speedups than POP by effectively breaking the problem into smaller and easier subproblems. Notably, DeDe and DeDe both reach a normalized proportional fairness of over 1 within 100 seconds. In contrast, POP-4 and POP-16 require 3,053 seconds and 682 seconds, respectively, to achieve comparable allocation quality, while POP-64 yields a significantly lower fairness score of only 0.31.
@@ 7.1.2 Traffic Engineering
Figure 6: Results of the traffic engineering problem variant that maximizes the total flow.
@@
We adopt the evaluation setup in Teal [65], with a focus on its largest topology—a 1,739-node network adapted from an internet topology for WAN purposes. The traffic matrices are collected from the production WAN of a global cloud provider and mapped onto the above topology.
@@
Maximize total flow.  Figure 6 compares DeDe to the baselines in maximizing total flow. DeDe achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, showing a superior trade-off between satisfied demand and computation time compared with Exact sol., POP, and Pinning. Its idealized variant, DeDe, achieves a satisfied demand that is 8% higher than DeDe within 10 seconds. While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer, at 1,658 seconds. POP-16 and POP-64 improve parallelism by dividing the workload into more subproblems, reducing computation times to 456 seconds and 380 seconds, respectively. Note that the speedup of POP-64 is limited compared with POP-16 as the number of cores per subproblem drops from 4 to 1. This increased parallelism also comes at a cost, with satisfied demand decreasing to 87.4% for POP-16 and 81.6% for POP-64. Pinning focuses on top demands to reduce problem size but remains constrained by the sequential nature of its optimization solvers. As a result, it achieves a modest speedup, requiring 149 seconds to reach 87.3% satisfied demand. Teal, a learning-accelerated approach, leverages massive GPU parallelism to achieve 89% satisfied demand in only 1 second. However, it relies on a highly customized machine-learning framework, demanding significant human effort for design and training.
@@
Figure 7: Results of the traffic engineering problem variant that minimizes the max link utilization.
Figure 8: Results of the load balancing problem where the objective is to minimize shard movements.
@@
Min-max link utilization.  Figure 7 compares DeDe with other approaches in minimizing maximum link utilization. Here, the link utilization metric serves as proxy for network congestion and is allowed to exceed 100% during optimization. DeDe attains a maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds. Exact sol. reaches 1.63 with a longer computation time of 35 seconds. This is faster than maximizing total flow because minimizing max link utilization is computationally less challenging and thus requires fewer iterations. Pinning takes 42.5 seconds, making it slower than DeDe and slightly slower than Exact sol., as it requires additional time to rebuild the problem when the top 10% of node pairs change. POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 within 9, 2, and 0.4 seconds on average, respectively. Although POP-16 is faster than DeDe, a fairer comparison between POP-16 and DeDe, both simulating parallelism, reveals that DeDe achieves a comparable computation time to POP-16. The domain-customized Teal takes only 0.3 seconds on a GPU to attain a maximum link utilization of 1.74.
7.1.3 Load Balancing
@@
We follow the load balancing settings in POP [44], except for increasing the scale to 2,048 data shards distributed across 256 servers. In each round, a new shard-to-server allocation is computed based on updated query loads. We simulate for 100 rounds, using the first 20 as a warm-up. To ensure the problem’s feasibility, we set the tolerance parameter $\epsilon$ to 0.1.
Minimize shard movements.  Figure 8 compares DeDe with baselines in minimizing shard movements. DeDe achieves an average of 20.1 shard movements in 15 seconds, better than POP-4’s 21.5 shard movements within 133 seconds. In faster scenarios, DeDe achieves 25.4 shard movements in 5 seconds, while POP-16 achieves an average movement of 26 in a slighter less time of 3.7 seconds. For a fair comparison, DeDe consistently achieves around 16% fewer shard movements than POP, given a similar time limit. Exact sol. requires 20.9 shard movements on average, but is significantly slower with an average of 4,820 seconds. This slower computation is due to the NP-hard nature of solving the mixed-integer linear program underlying load balancing. Notably, Exact sol. does not produce the minimum possible shard movements because optimizing each scheduling round independently does not guarantee an optimal solution across the entire series. At the opposite end of the time-accuracy trade-off, the greedy heuristic provides a rapid but less accurate solution, completing in just 2 milliseconds but averaging 73 shard movements after naively fixing its constraint violations.
@@
(c) Comparing different optimization methods.
@@ 7.2 Robustness
Using traffic engineering as an example, we evaluate the robustness of DeDe and other baselines against changes in problem granularity, temporal dynamics, spatial redistribution, and link failures. In these experiments, DeDe is run for 30 seconds, and we normalize the satisfied traffic demand by the optimal one achieved by Exact sol.
Changes of problem granularity.  The granular problem structure required by POP, where each demand only requests a small fraction of interchangeable resources, can be impractical. In traffic engineering, for example, demands are constrained by the link capacity on specific pre-configured paths. Similarly, in cluster scheduling, 33% of GPU tasks in production clusters are limited to specific GPU types [59]. In Figure 9(a), we present results on varying levels of problem granularity using traffic engineering as an example. To quantify resource interchangeability, we use the mean edge betweenness centrality [9, 10], which measures the average percentage of demands served by a given edge.
@@
Temporal dynamics.  Next, we introduce temporal fluctuations [65] to the traffic matrices for robustness assessment. For each demand, we calculate the variance $\sigma^{2}$ in its changes between consecutive time slots and create a new normal distribution $N(0,k\sigma^{2})$. We choose the values of $k$ to be 2, 5, 10, and 20. Next, we randomly draw a sample from the newly created normal distribution and add it to each demand in every time slot, simulating temporal fluctuations.
@@
Spatial redistribution.  To evaluate robustness against spatial redistribution, we reassign traffic demands among node pairs. Specifically, we adjust the top 10% of demands, which initially account for 88.4% of the total volume, so that they take up 80%, 60%, 40%, and 20% of the total volume instead. Figure 9(c) shows that DeDe consistently fulfills the highest demand across all spatial distributions. Pinning’s performance decreases by 8.3%, as its strategy of focusing only on top demands relies on the presence of a heavy-tailed demand distribution. Teal has a slight decrease of up to 2.4% due to unseen demand distribution in training. This is a relatively smaller decrease than the temporal fluctuation, because Teal can still handle simple, evenly distributed traffic demands. POP has a slight increase of 1.4% as demands are more granular, where no demand requires a significant fraction of the total resources.
Link failures.  Widespread failures across inter-datacenter links are relatively rare in real-world operations unless triggered by physical fiber cuts [52]. However, we simulate severe failure scenarios, introducing 50, 100, and 200 simultaneous link failures. As shown in Figure 11, the satisfied demand declines consistently across all methods. Specifically, all approaches drop by 0.6–1.1% under 50 link failures, 1.5–2.1% under 100 link failures, and 3–4.2% under 200 link failures. These drops stem from the fact that the failed links represent only a small portion of the total 8,558 links, allowing all approaches to recompute flow allocation given adequate time.
@@ 7.3 Micro-benchmarks
We conduct micro-benchmarks to evaluate the impact of DeDe’s design components, with all experiments performed in maximizing total flow in traffic engineering.
@@
Speedup when varying CPU cores.  Figure 10(a) evaluates the speedup of DeDe, DeDe, and Exact sol. with 1, 4, 16, and 64 CPU cores. DeDe achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores. DeDe also achieves an almost linear speedup when the number of CPU cores grows from 1 to 16. However, this linearity diminishes with 64 cores, where DeDe only achieves 18.2$\times$ speedup. This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in §7.1.1. In contrast to the linear speedup of DeDe and DeDe, the speedup of Exact sol. is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores. This is because of commercial solvers’ sequential nature, e.g., the simplex method [46] in linear programming requires thousands to millions of small steps to reach the optimal solution.
Convergence rate.  Teal generates a coarse, initial solution to traffic engineering using neural networks and fine-tunes it through ADMM. While Teal also “decouples” constraints, it does not further “decompose” the problem as DeDe does. Instead, it analytically and approximately minimizes a problem-specific augmented Lagrangian, directly projecting the result onto the non-negative feasible region. As a specialized method, Teal converges rapidly, achieving 90% satisfied demand within 10 seconds, as shown in Figure 10(b). In comparison, DeDe requires $\sim$30 seconds to attain comparable performance. In each optimization interval, DeDe is warm-started using the previous solution by default. When initialized instead with Teal’s neural network output, DeDe converges at a similar rate. In contrast, employing a naive initialization—where each demand is equally split across available paths—reduces DeDe’s convergence speed by nearly half.
Alternative optimization methods.  Figure 10(c) evaluates alternative constrained optimization techniques for maximizing total flow in traffic engineering—the penalty method [4] and the augmented Lagrangian method [23], comparing them with the ADMM-based approach used by DeDe. Both methods aim to solve the reformulated optimization problem in Equation 3.1 by jointly (and naively) optimizing variables $x$ and $z$, but they differ in how they enforce constraints. The penalty method introduces penalty terms into the objective function to discourage constraint violations. It starts with a small penalty coefficient and gradually increases it toward infinity, thereby driving the solution toward feasibility. However, this process typically involves solving a series of increasingly ill-conditioned problems, resulting in slow convergence. In our evaluation, it was more than 30$\times$ slower than DeDe in achieving a satisfied demand of over 90%. The augmented Lagrangian method improves upon the penalty method by combining both penalty terms and Lagrange multipliers. This helps enhance convergence stability and typically reduces the number of required iterations. However, unlike the ADMM approach used in DeDe, it lacks the benefit of decomposition and parallelism. As a result, despite its faster convergence over the basic penalty method, it remains over 3$\times$ slower than DeDe in reaching a satisfied demand of over 90%.